# Federated Learning for Traffic Flow Prediction with Synthetic Data Augmentation

### v1

@@ 6.1. Trajectory Data Generation
FedOpt 0.2330 23.68 0.1287 22.22 0.1236 19.81 0.1193 18.93 0.1120 17.78 0.1181 18.72
FedProx 0.2337 24.02 0.1283 23.44 0.1294 20.75 0.1217 19.89 0.1194 19.66 0.1183 18.30
FedTPS 0.2308 23.27 0.1229 20.93 0.1184 18.89 0.1148 18.65 0.1096 17.21 0.1144 18.34
4 Non-FL 0.2613 26.75 0.1731 27.43 0.1760 25.51 0.1699 25.23 0.1527 21.80 0.1468 20.86
@@
FedOpt 0.2466 23.92 0.1502 23.24 0.1482 21.84 0.1419 21.78 0.1373 20.59 0.1422 20.42
FedProx 0.2474 24.21 0.1537 24.68 0.1489 22.06 0.1455 21.23 0.1419 21.44 0.1428 20.95
FedTPS 0.2470 24.00 0.1487 23.13 0.1434 21.41 0.1405 21.19 0.1428 20.46 0.1405 20.70
6 Non-FL 0.2719 27.08 0.1868 28.32 0.1928 26.03 0.1851 26.03 0.1825 24.21 0.1660 22.35
@@
FedOpt 0.2611 25.30 0.1691 24.89 0.1685 23.52 0.1620 22.50 0.1602 21.81 0.1622 22.30
FedProx 0.2628 25.47 0.1714 25.47 0.1680 23.62 0.1645 22.71 0.1668 22.84 0.1645 22.33
FedTPS 0.2611 25.08 0.1654 24.44 0.1637 22.89 0.1600 22.35 0.1668 22.35 0.1591 21.78
Table 2 shows that the data heterogeneity in the partitioned datasets increases in the settings with more client partitions. This can cause the global model to struggle generalising across the multiple data silos, which leads to model performance degradation with higher client numbers. Furthermore, since the client’s partitioned datasets decrease in size with more clients, this introduces statistical challenges for the diffusion model stemming from smaller local datasets, which can slow convergence for the global model. Yet the samples generated throughout all client cases demonstrate high degrees of similarity with the real data, as the structural similarity index is very close to 1 and the nRMSE remains low throughout the different cases.
Figure 6. Visualisation of generated trajectories for the city of Chengdu from federated DiffTraj model over varying client cases.
@@
Generated trajectory samples Chengdu
6.2. Traffic Flow Prediction
Table 4 presents the results of the traffic prediction models, trained under different FL paradigms with varying client numbers, with respect to normalised Mean Absolute Error (nMAE) and Mean Absolute Percentage Error (MAPE) metrics. In general, FedTPS consistently outperforms other FL frameworks for the Chengdu dataset. Our reasoning for FedTPS’s improved performance relative to other FL frameworks is due to the increased amount of data available to the clients during training, as well as the reduction of data heterogeneity stemming from the shared synthetic dataset. Through synthetic data augmentation, we ensure that each client’s dataset is more representative of the overall data distribution, and increase the amount of information available to each client. By reducing the degree of data heterogeneity the impact of non-IIDness on client training which may manifest as noise in the gradient updates is ameliorated. This improves the stability of the training by improving the parameter fusion process, and consequently the global model performance.
Comparing FedAvg and FedOpt, there is no noticeable difference between the two mechanisms as FedAvg outperforms FedOpt in 6 cases, while FedOpt outperforms FedAvg in 7 cases. This is consistent with wider research on cross-silo FL with real-world data, which found that no State-of-the-Art FL framework consistently outperformed FedAvg (Li et al., 2022). FedProx generally performs worse than the other FL frameworks. It should be noted that the data augmentation framework of FedTPS can be used in conjunction to other FL framework which optimise the aggregation protocol for global model updates or the client selection strategy, such as any of the FL baselines presented in this work.
With regards to the traffic flow prediction models, TAU consistently outperforms other State-of-the-Art models. When trained with FedTPS, GATAU yields the best results under a six client setting, and has the lower nMAE in a four client setting. The Graph Attention mechanism has more trainable parameters than the convolution operations employed by TAU in its encoder-decoder structure. Therefore data augmentation may be more impactful for this model. At lower client numbers, the benefits of FedTPS are more noticeable. This is likely due to the performance of the trajectory generation model improving at low client numbers, hence leading to higher quality synthetic trajectories generated. Furthermore, as shown in Table 2, data heterogeneity increases when partitioning the data for higher client number cases, and thus, the optimisation objective in local training for a client varies more. This impacts the global model developed in a federated setting, which may result in worse performing global models.
6.3. Effects of Synthetic Pre-training on Global Model Training
Figure 7 details the training loss and evaluation metrics of select traffic prediction models under FL settings, with and without pre-training. With pre-training, the models can be seen to converge more rapidly than the FL settings without pre-training. However, by global round 60, the models will have generally converged to the same training loss. Hence, while pre-training with synthetic data improves the convergence speed of the global model noticeably, it offers no improved model performance. The random initialisation of model parameters causes the training loss and evaluation metrics at global round 1 to be very high for models without pre-training. However, with synthetic data pre-training, the starting point of the model is competitive, as can be seen from the low evaluation metric scores at global round 1.
Figure 7. Effects of Pre-Training models with Synthetic Data on model training process.
@@ 6.3. Effects of Synthetic Pre-training on Global Model Training
Training losses for models with and without pre-training
6.4. Synthetic Data Augmentation Single-Client Setting
Table 5. Comparison of various Spatio-Temporal model performances for prediction of traffic at 6 time-steps in future. Assessed using nMAE and MAPE metrics, when there is a single client.
GRU STGCN DCRNN GWNET TAU GATAU
City Augmented Data nMAE MAPE nMAE MAPE nMAE MAPE nMAE MAPE nMAE MAPE nMAE MAPE
Chengdu No 0.2419 34.47 0.1409 28.58 0.1379 23.33 0.1320 21.95 0.1003 17.63 0.1039 16.83
Yes 0.2338 29.74 0.1323 27.89 0.1263 22.01 0.1247 20.68 0.0928 16.02 0.0952 16.21
Table 5 details the effects of synthetic data augmentation for multiple traffic flow prediction models in a single client case. For this dataset, it is clear in all cases that synthetic data augmentation can significantly improve model performance. In a single client case, where there is limited traffic regional inflow data, adding synthetic data can increase the amount of information available during training. Furthermore, a single client case would also yield a diffusion model which generates the highest fidelity trajectories; it can be seen from Table 3 that increasing the number of clients leads to a degradation in the quality of the generated trajectories.
6.5. Effect of Client Rounds and Server Rounds
@@
Table 6. Comparison of selected traffic flow prediction models trained in a federated setting employing FedAvg and FedTPS as the ratio of local and global training rounds are varied.
@@
2 STGCN FedAvg 10 0.1301 23.22
20 0.1327 24.27
FedTPS 10 0.1229 21.84
20 0.1228 22.03
GWNET FedAvg 10 0.1165 19.00
20 0.1238 21.10
FedTPS 10 0.1156 19.19
20 0.1149 19.40
GATAU FedAvg 10 0.1183 18.42
20 0.1230 18.62
FedTPS 10 0.1179 17.93
20 0.1169 17.88
4 STGCN FedAvg 10 0.1573 24.93
20 0.1564 24.86
FedTPS 10 0.1489 23.54
20 0.1534 24.45
GWNET FedAvg 10 0.1442 21.44
20 0.1474 21.84
FedTPS 10 0.1401 21.27
20 0.1398 21.08
GATAU FedAvg 10 0.1434 20.67
20 0.1417 20.50
FedTPS 10 0.1412 21.01
20 0.1395 20.20
6 STGCN FedAvg 10 0.1744 26.18
20 0.1752 26.25
FedTPS 10 0.1674 24.65
20 0.1735 25.93
GWNET FedAvg 10 0.1643 22.70
20 0.1646 23.15
FedTPS 10 0.1608 22.39
20 0.1622 22.48
GATAU FedAvg 10 0.1627 22.53
20 0.1628 21.95
FedTPS 10 0.1601 21.86
20 0.1595 21.84
@@
Table 6 illustrates the results of this study, which explores varying the number of local training rounds per global aggregation step. FedTPS consistently demonstrates lowest error metrics than FedAvg. The decrease in data heterogeneity stemming from the shared synthetic dataset, coupled with the increased information available to each client, means that even in cases where there are greater local training rounds relative to global aggregation rounds, the global model convergence is enhanced. This may be because the synthetic data reduces the variance between client updates and promotes more stable and consistent parameter updates. Thus, FedTPS improves the robustness of the federated learning process against the challenges posed by non-IID data distributions and limited communication rounds.
6.6. Communication Costs
Table 7 presents FedTPS’ communication cost for the distribution of the synthetic data to each federated client in. as well as the communication cost of the model parameters which occurs when the server distributes the global model to clients.
FedTPS distributes the generated synthetic data to clients only at the initial round. Compared to the communication cost for a single round of the global model distribution process, FedTPS increases communication cost from 44.7% for the STGCN model, to 0.8% for the TAU model. The global model distribution process is a reoccurring communication cost, and hence over the entire federated training process the added communication cost of FedTPS becomes much less significant.
Table 7 also demonstrates that the GATAU model incurs less communication costs than the TAU model. It should be noted that for cross-silo FL settings, communication bottlenecks are not as prevalent; server-to-client connections represent high-bandwidth data center network connections. Furthermore, complete client participation is important for cross-silo settings as contributions from every client are critical.
Table 7. Communication costs of FedAvg and FedTPS for selected traffic flow prediction models for different client numbers, in MB.
Initial Comm. Cost per Client Comm. Cost per Client per FL Round
City FedTPS STGCN GWNET TAU GATAU
Chengdu 0.38 0.85 1.21 47.83 38.70
7. Conclusion
@@
This paper addresses the task of Federated Learning for traffic flow prediction applications, and presents a new framework referred to as FedTPS which facilitates synthetic data augmentation of client datasets. By reducing data heterogeneity among clients and enhancing the information available during local training, FedTPS is able to consistently outperform other FL mechanisms for a variety of different client settings and different traffic flow prediction models. An FL setting to traffic flow prediction tasks is emulated using a large-scale ride sharing dataset for the city of Chengdu. The trajectory data is partitioned into subsets to emulate individual organisations which collect their own data. Organisations can then collaborate to train models in a federated manner, without infringing upon data privacy. FedTPS first trains a federated synthetic data generation model, and then generates a synthetic regional traffic flow dataset from the learned global distribution of trajectories. This synthetic data is then distributed to all clients prior to the traffic flow prediction task so they may augment their local datasets. Future work could apply FedTPS in highly non-IID data settings.

### v2

@@ VI-A Traffic Flow Data Generation
FedOpt 0.2190 23.90 0.1186 21.88 0.1152 20.26 0.1149 20.19 0.1043 17.91 0.1144 19.01
FedProx 0.2167 25.02 0.1215 23.07 0.1214 21.10 0.1186 21.28 0.1122 18.61 0.1129 18.88
FedTPS 0.2143 22.91 0.1122 19.88 0.1065 17.96 0.1080 18.69 0.1019 17.11 0.1090 18.02
4 Non-FL 0.2512 27.65 0.1758 30.47 0.1750 26.91 0.1651 25.01 0.1487 22.66 0.1472 21.59
@@
FedOpt 0.2329 24.33 0.1442 24.26 0.1426 22.94 0.1384 22.02 0.1289 20.44 0.1373 21.22
FedProx 0.2325 24.72 0.1455 24.75 0.1430 22.93 0.1386 22.38 0.1341 21.36 0.1373 21.21
FedTPS 0.2272 23.28 0.1360 22.62 0.1316 21.22 0.1282 20.39 0.1335 20.89 0.1299 20.38
6 Non-FL 0.2614 32.31 0.1962 33.88 0.1954 31.03 0.1877 28.70 0.1760 26.66 0.1660 26.01
@@
FedOpt 0.2459 25.35 0.1602 24.99 0.1614 24.01 0.1554 23.02 0.1488 21.91 0.1562 22.48
FedProx 0.2442 25.65 0.1614 25.45 0.1620 24.23 0.1557 23.39 0.1528 22.57 0.1557 22.66
FedTPS 0.2402 24.57 0.1551 24.09 0.1525 22.62 0.1502 22.15 0.1542 22.69 0.1502 22.05
8 Non-FL 0.2738 28.08 0.2091 31.36 0.2114 29.24 0.2057 28.62 0.2038 27.52 0.1836 25.78
@@
FedOpt 0.2586 27.23 0.1805 27.14 0.1813 25.62 0.1756 25.64 0.1687 23.79 0.1760 24.47
FedProx 0.2597 27.18 0.1855 27.81 0.1832 25.99 0.1763 25.34 0.1706 24.23 0.1782 24.74
FedTPS 0.2556 26.83 0.1767 26.50 0.1741 24.60 0.1691 24.39 0.1801 24.87 0.1725 23.79
Xi’an 2 Non-FL 0.2389 44.62 0.1719 30.44 0.1716 28.74 0.1516 24.91 0.1387 23.25 0.1350 21.21
@@
FedOpt 0.2133 33.44 0.1369 20.13 0.1351 20.40 0.1395 21.51 0.1276 18.81 0.1276 20.18
FedProx 0.2151 33.31 0.1399 21.56 0.1352 21.29 0.1376 21.63 0.1324 20.35 0.1300 21.05
FedTPS 0.2096 32.09 0.1256 18.23 0.1284 19.59 0.1302 19.96 0.1304 18.60 0.1240 19.35
4 Non-FL 0.2587 38.18 0.2173 34.52 0.2071 32.34 0.1871 28.58 0.1852 26.99 0.1657 25.11
[…]
FedOpt 0.2354 27.47 0.1698 22.95 0.1682 23.53 0.1632 23.41 0.1627 22.32 0.1597 22.40
FedProx 0.2362 27.17 0.1736 23.93 0.1693 23.39 0.1632 23.53 0.1618 22.95 0.1594 22.43
FedTPS 0.2345 26.68 0.1621 22.06 0.1613 22.94 0.1569 22.64 0.1709 23.37 0.1586 21.81
6 Non-FL 0.2819 33.64 0.2230 32.15 0.2271 32.56 0.2099 30.39 0.2271 31.77 0.1913 26.23
[…]
FedOpt 0.2535 26.52 0.1938 25.96 0.1975 25.69 0.1930 26.82 0.1872 24.25 0.1819 24.54
FedProx 0.2559 26.69 0.1983 26.85 0.1983 26.08 0.1959 26.49 0.1880 25.05 0.1839 24.83
FedTPS 0.2510 26.26 0.1893 25.79 0.1913 25.39 0.1913 26.34 0.2020 25.91 0.1827 24.60
8 Non-FL 0.2935 28.92 0.2403 31.80 0.2480 30.99 0.2359 30.10 0.2529 31.66 0.2238 27.41
[…]
FedOpt 0.2694 27.79 0.2129 28.20 0.2173 28.01 0.2085 27.92 0.2090 26.52 0.2052 26.70
FedProx 0.2683 28.02 0.2140 28.63 0.2184 28.40 0.2101 27.94 0.2096 26.99 0.2057 26.71
FedTPS 0.2677 27.74 0.2107 28.10 0.2129 27.68 0.2074 27.70 0.2255 28.02 0.2057 26.31
Table II shows that the data heterogeneity in the partitioned datasets increases when there are more clients in the federated setting. This can cause the global model to struggle generalising across the multiple data silos, which leads to model performance degradation with higher client numbers, and slower global model convergence. Yet the samples generated throughout all client cases demonstrate high degrees of similarity with the real data, as the structural similarity index and temporal correlation among regions remain close to 1, and the nRMSE remains low throughout the different cases.
VI-B Traffic Flow Prediction
Table V presents the results of the traffic prediction models, trained under different FL paradigms with varying client numbers, with respect to nMAE and MAPE metrics. FedTPS consistently outperforms other FL frameworks for the Chengdu dataset. Our reasoning for FedTPS’s improved performance relative to other FL frameworks is due to the increased amount of data available to the clients during training, as well as the data heterogeneity reduction stemming from the shared synthetic dataset. Through synthetic data augmentation, we ensure that each client’s dataset is more representative of the overall data distribution, and increase the amount of information available to each client.
Comparing FedAvg and FedOpt, there is no noticeable difference between the two mechanisms. This is consistent with wider research on cross-silo FL with real-world data, which found that no State-of-the-Art FL framework consistently outperformed FedAvg [39]. FedProx generally performs worse than the other FL frameworks. It should be noted that the data augmentation framework of FedTPS can be used in conjunction to other FL framework which optimise the aggregation protocol for global model updates or client selection strategy, such as any of the FL baselines presented in this work.
With regards to the traffic flow prediction models, TAU consistently outperforms other State-of-the-Art models. When trained with FedTPS, GATAU can yield better results, as shown in all the client cases for the Xi’an dataset. The Graph Attention mechanism can learn cross-region relationships from the entire city, whereas the convolutional operations implemented in the TAU restrict spatial dependencies modelling to neighbouring regions. Therefore, data augmentation may benefit GATAU more, as there is more data available to learn cross-region dependencies.
Aside from the two client case for both cities, training the TAU model under the FedTPS framework degrades the performance of the global model relative to FedAvg. This may stem from the fact that the TAU model mainly focuses on capturing temporal dependencies of the data, and Table IV shows that TC decreases more significantly than SSI in higher client cases.
At lower client numbers, the benefits of FedTPS are more noticeable. This is likely due to the performance of the trajectory generation model improving at low client numbers, hence leading to higher quality synthetic trajectories generated. Furthermore, as shown in Table II, data heterogeneity increases when partitioning the data for higher client number cases, and thus, the optimisation objective in local training for a client varies more. This impacts the federated global model, which may result in worse performing global models.
VI-C Communication Costs
Table VI presents the communication costs for the distribution of the global model to each client in the federated setting, for TFDiff and the traffic prediction models. The generative TFDiff model which is trained in the first stage of FedTPS requires less communication cost than TAU or GATAU. Table VI also demonstrates that the GATAU model incurs less communication costs than the TAU model. It should be noted that for cross-silo FL settings, communication bottlenecks are not as prevalent; server-to-client connections represent high-bandwidth data center network connections. Furthermore, complete client participation is important for cross-silo settings as contributions from every client are critical.
TABLE VI: Communication costs of FedAvg and FedTPS for selected traffic flow prediction models for different client numbers, in MB.
Comm. Cost per Client per FL Round
TFDiff GRU STGCN DCRNN GWNET TAU GATAU
47.52 0.27 0.87 1.79 1.24 50.76 48.40
VI-D Effects of Synthetic Pre-training on Global Model Training
Figure 7 details the training loss and evaluation metrics of select traffic prediction models under FL settings, with and without pre-training. With pre-training, the models converge more rapidly than the FL settings without pre-training. However, by global round 60, the models will have generally converged to the same training loss. Hence, while pre-training with synthetic data improves the convergence speed of the global model noticeably, it offers no improved model performance. The random initialisation of model parameters causes the training loss and evaluation metrics at global round 1 to be very high for models without pre-training. However, with pre-training, the starting point of the model is competitive, as can be seen from the low metric scores at global round 1.
Figure 7: Effects of Pre-Training models with Synthetic Data on model training process.
VI-E Effect of Client Rounds and Server Rounds
@@
TABLE VII: Comparison of selected traffic flow prediction models trained in a federated setting employing FedAvg and FedTPS as the ratio of local and global training rounds are varied.
@@
2 STGCN FedAvg 10 0.1213 23.10
20 0.1210 23.23
FedTPS 10 0.1153 20.12
20 0.1145 20.58
GWNET FedAvg 10 0.1113 18.88
20 0.1163 19.70
FedTPS 10 0.1085 17.95
20 0.1097 19.22
GATAU FedAvg 10 0.1192 20.46
20 0.1124 20.08
FedTPS 10 0.1101 17.93
20 0.1107 18.21
4 STGCN FedAvg 10 0.1466 23.76
20 0.1480 24.58
FedTPS 10 0.1410 23.68
20 0.1414 23.39
GWNET FedAvg 10 0.1384 21.85
20 0.1383 21.79
FedTPS 10 0.1277 21.20
20 0.1308 21.23
GATAU FedAvg 10 0.1374 20.98
20 0.1358 21.17
FedTPS 10 0.1322 20.98
20 0.1329 20.39
6 STGCN FedAvg 10 0.1633 25.95
20 0.1637 25.76
FedTPS 10 0.1572 24.45
20 0.1579 24.53
GWNET FedAvg 10 0.1563 23.04
20 0.1582 23.58
FedTPS 10 0.1502 22.26
20 0.1525 22.24
GATAU FedAvg 10 0.1559 22.33
20 0.1562 22.51
FedTPS 10 0.1516 22.04
20 0.1499 21.80
8 STGCN FedAvg 10 0.1860 28.16
20 0.1843 27.89
FedTPS 10 0.1801 26.91
20 0.1854 27.73
GWNET FedAvg 10 0.1764 25.01
20 0.1791 25.52
FedTPS 10 0.1710 24.57
20 0.1708 24.43
GATAU FedAvg 10 0.1772 24.50
20 0.1775 24.14
FedTPS 10 0.1715 23.96
20 0.1734 23.75
@@
Table VII illustrates the results of this study, which explores varying the number of local training rounds per global aggregation step. FedTPS consistently demonstrates lowest error metrics than FedAvg. The decrease in data heterogeneity stemming from the shared synthetic dataset, coupled with the increased information available to each client, means that even in cases where there are greater local training rounds relative to global aggregation rounds, the global model performance is enhanced. This may be because the synthetic data reduces the variance between client updates and promotes more stable and consistent parameter updates. Thus, FedTPS improves the robustness of the FL process against the challenges posed by non-IID data distributions and limited communication rounds.
VII Conclusion
@@
This paper addresses the task of Federated Learning for traffic flow prediction applications, and presents a new framework referred to as FedTPS which facilitates synthetic data augmentation of client datasets. By reducing data heterogeneity among clients and enhancing the information available during local training, FedTPS is able to consistently outperform other FL mechanisms for a variety of different client settings and different traffic flow prediction models. An FL setting to traffic flow prediction tasks is emulated using a large-scale ride sharing dataset for the cities of Chengdu and Xi’an. The trajectory data is partitioned into subsets to emulate individual organisations which collect their own data. Organisations can then collaborate to train models in a federated manner, without infringing upon data privacy. FedTPS first trains a federated synthetic data generation model, TFDiff, and then generates a synthetic regional traffic flow dataset from the learned global distribution of trajectories. This synthetic data is then distributed to all clients prior to the traffic flow prediction task so they may augment their local datasets. Future work could apply FedTPS in highly non-IID data settings.